# Generalization of terms via universal algebra

### v1

@@ 4.3. A universal-algebraic methodology for the e-generalization type
Assume first that the type of $\mathscr{G}(h)$ is nullary or infinitary; then we prove that there cannot be a finite minimal complete set of solutions to $h$. Assume by way of contradiction that there is, and call it =$\mathscr{M}=\{g_{1},\ldots,g_{n}\}$, where each $g_{i}:{\bf F}_{\mathsf{V}}(z)\to{\textbf{P}}_{i}$ for some projective algebra ${\textbf{P}}_{i}\in\mathsf{V}$. Consider now the set of congruences $\{\theta_{i}=\ker(g_{i}):i=1,\ldots,n\}\subseteq\mathscr{G}(h)$; since the type of $\mathscr{G}(h)$ is nullary or infinitary, there must be an exact congruence $\theta\in\mathscr{G}(h)$ such that $\theta\not\subseteq\theta_{i}$ for all $i=1,\ldots,n$. For otherwise one would have that the type of $\mathscr{G}(h)$ is finitary.
@@
For (2), suppose now that the type of $\mathscr{G}(h)$ is finitary. Then the same argument as above, setting $n=1$ and =$\mathscr{M}=\{g_{1}\}$, shows that the e-generalization type of $h$ cannot be unitary. ∎
@@
Example 4.22.
In Example 4.14 we have seen that every variety $\mathsf{V}$ whose 1-generated free algebra is trivial has the 1ESP property. This includes lattices, semilattices, and more generally all varieties without constants and whose operations are idempotent. Since the 1-generated free algebra is trivial, the type of $\mathscr{G}(h)$ is clearly unitary for all problems $h$; it follows from Corollary 4.20 that every such variety has unitary $\kappa$-type for every cardinal $\kappa\leq\omega$. $\Box$
@@
It is proved in [16] that idempotent anti-unification is infinitary, which seems to contradict the result outlined in the above example. However, the result in [16] makes key use of two constants, assumed to be in the language, in order to show that the type is (at least) infinitary. If we add constants, the 1-generated free algebra is not trivial, and therefore there is no contradiction. We also note that our result is incomparable with the one in [16], in the sense that while ours is more general with respect to the fact that we do not restrict the equational theory to idempotency (our result holds for any equational theory that includes idempotency of all symbols), we do need to restrict to not having constants.
@@
The previous theorem entails a sufficient condition for varieties to have unitary e-generalization type.
Corollary 4.25.
@@
If $h:{\bf F}_{\mathsf{V}}(z)\to\prod_{k=1}^{m}{\textbf{E}}_{k}$ is any e-generalization problem, under our hypothesis, $\prod_{k=1}^{m}{\textbf{E}}_{k}$ is finitely generated and projective and hence its type is unitary by Theorem 4.24. Thus, the $\kappa$-type of $\mathsf{V}$, for all cardinals $\kappa\leq\omega$, is unitary as well. ∎
Example 4.26.
Abelian groups, commutative monoids, and commutative semigroups have unitary e-generalization type. Indeed, in Example 4.8 we have seen that the only 1-generated exact algebras are the free 1-generated algebra itself, ${\bf F}_{\mathsf{V}}(z)$ and, in all cases except commutative semigroups, the trivial algebra (which in such cases is projective). Thus, all nontrivial finite direct products of 1-generated exact algebras are finite powers of ${\bf F}_{\mathsf{V}}(z)$.
In all three cases, the $n$-th power ${\bf F}_{\mathsf{V}}(z)^{n}$ is isomorphic to the $n$-generated free algebra. To see this, notice that for all three varieties under consideration the $n$-generated free algebra is the algebra of polynomials (over the appropriate language) written over $n$-variables. Specifically, for abelian groups, commutative monoids, and commutative semigroups one gets $\mathbb{Z}[x_{1},\ldots,x_{n}]$, $\mathbb{N}[x_{1},\ldots,x_{n}]$, and $\mathbb{N}^{+}[x_{1},\ldots,x_{n}]$ respectively. It is then easy to realize that these are isomorphic to, respectively, ${\textbf{Z}}^{n}$, ${\textbf{N}}^{n}$, and $({\textbf{N}}^{+})^{n}$. Therefore, finite direct products of 1-generated exact algebras are (free hence) projective, and we can apply Corollary 4.25 and conclude that all the three varieties above have unitary e-generalization type. [^2] Baader in [5] studies the generalization type of what are called commutative theories, which encompass commutative monoids and abelian groups, showing indeed unitary type. However, the generalization order considered in that paper is seemingly different from the one we consider here, as it involves ordering the testifying substitutions and not just the terms.
Note that if we drop commutativity and consider groups, monoids, and semigroups, finite direct products of 1-generated exact algebras are not projective. Indeed, the 1-generated exact algebras are the same as their commutative versions (see Example 4.8) and hence so are their direct products, which are then commutative. However, these direct products cannot embed into any free algebra since no nontrivial subalgebra of free groups/monoids/semigroups is commutative. Hence finite direct products of 1-generated exact algebras are neither exact nor projective. $\Box$
@@ 5. Applications to (algebraizable) logics
Given this general perspective, we now investigate some of the most well-known (algebraizable) logics and obtain examples of unitary type for the varieties of Boolean, Gödel, and Kleene algebras.
@@ 5.1. Classical logic
Figure 2. ${\bf F}_{\mathsf{BA}}(z)$ and its congruence lattice, where $\langle z,1\rangle$ and $\langle\neg z,1\rangle$ represent, respectively, the congruences generated by $(z,1)$ and ($(\neg z,1)$.
@@
Observe that one could have obtained the above result also utilizing Corollary 4.25, since all finitely generated Boolean algebras, and hence in particular finite direct products of 1-generated exact Boolean algebras, are projective.
5.2. Gödel-Dummett logic
@@
Since all non-trivial finite Gödel algebras are projective (see [24, Proposition 2.5] for a proof based on a Priestley-like duality, and [1, Theorem 5.17] for an algebraic proof), $\mathsf{GA}$ has the 1EP property. However, it is not hard to see that the 1ESP fails, as the following example shows.
@@
Nonetheless, we can use Corollary 4.25 to prove that Gödel algebras have unitary e-generalization type. Indeed, all non-trivial finite Gödel algebras are projective, and since Gödel algebras are a locally finite variety, finite direct products of 1-generated exact algebras are finite. We can then conclude the following.
@@
Let us finally consider the 3-valued Kleene logic, also known as Kleene’s “strong logic of indeterminacy”; this is a generalization of classical logic which adds to the intended model of the logic a third indeterminate value to the truth and falsum constants [38]. Its algebraic semantics, the variety $\mathsf{KA}$ of Kleene algebras, is a subvariety of bounded distributive lattices with an involution $\neg$ [37]. In particular $\mathsf{KA}$ is axiomatized by $x\approx\neg\neg x$, De Morgan identity $x\wedge y\approx\neg(\neg x\lor\neg y)$, and Kleene identity $x\wedge\neg x\leq y\vee\neg y$. As it is well-known $\mathsf{KA}$ is generated by its standard model which is the 3-element algebra ${\textbf{K}}_{3}=(\{0,a,1\},\wedge,\vee,\neg,0,1)$ where $10<a<$ 1 and $\neg 0=1,\neg 1=0,\neg a=a$.
The free $1$-generated algebra ${\bf F}_{\mathsf{KA}}(z)$ is depicted in Figure 3, together with its quotients ${\bf 4}_{\mathsf{KA}}$ and ${\bf 2}_{\mathsf{KA}}$ (respectively, the 4-element and 2-element Boolean algebras seen as Kleene algebras), the Kleene chains ${\bf K}_{4}$ and ${\bf K}_{3}$, of four and three elements, and the trivial algebra.
00{0}${z\wedge\neg z}$${\neg z}$${z}$${z\vee\neg z}$11{1}1
0011{1}1
0011{1}1
0011{1}1
0011{1}1
@@
(4) $\neg x\leq x,x\land\neg y\leq\neg x\lor y\mbox{ implies }\neg y\leq y.$
We observe that if a Kleene algebra has a negation fixpoint $f=\neg f$ it is not exact, because the above quasi-equation (4) fails. Indeed, $\neg f\leq f$ and $f=f\wedge\neg 0\leq\neg f\vee 0=\neg f$ but $\neg 0=1>0$.
@@
The 1-generated exact Kleene algebras are ${\bf F}_{\mathsf{KA}}(1)$, ${\bf K}_{4}$, and ${\bf 2}_{\mathsf{KA}}$, and they are all projective.
@@
Let us show that, among the 1-generated algebras depicted in Figure 3, only ${\bf F}_{\mathsf{KA}}(1)$, ${\bf 4}_{\mathsf{KA}}$, ${\bf K}_{4}$, ${\bf 2}_{\mathsf{KA}}$ are exact. First, ${\bf F}_{\mathsf{KA}}(1)$ is a retract of itself and it is hence exact and projective. It is also clear that ${\bf 2}_{\mathsf{KA}}$ is a retract of ${\bf F}_{\mathsf{KA}}(z)$ and therefore it is projective and hence exact.
The same argument holds for ${\bf K}_{4}$. Indeed, naming its lattice reduct $10<\neg x<x<$ 1, consider the embedding into ${\bf F}_{\mathsf{KA}}(z)$ via the map $i$ such that $i(1)=1$, $i(x)=z\vee\neg z$, $i(\neg x)=z\wedge\neg z$, $i(0)=0$. The surjective homomorphism $j:{\bf F}_{\mathsf{KA}}(z)\to{\bf K}_{4}$ defined by $j(z)=j(z\lor\neg z)=x$ and $j(\neg z)=j(z\land\neg z)=\neg x$, composed with $i$ gives the identity on ${\bf 4}_{\mathsf{KA}}$, which is therefore a retract of the free algebra hence projective.
It is left to prove that the other 1-generated algebras are not exact. The trivial algebra is obviously not isomorphic to a subalgebra of any non-trivial algebra and therefore it cannot be exact. As for ${\textbf{4}}_{\mathsf{KA}}$, $1$ is not join irreducible and therefore the claim follows by Lemma 5.4. Finally, ${{\textbf{3}}}_{\mathsf{KA}}$ is not exact since it has a negation fixpoint and, as discussed above, this implies a failure of identity (4) which is necessary for exactness by Lemma 5.4. ∎
@@
(5) $\iota(x)=\bigwedge A-\{\neg a:x\leq_{{\textbf{A}}}a\}.$
@@
$\Delta$${\langle z\vee\neg z,1\rangle}$${\langle z,z\vee\neg z\rangle}$${\langle z,z\wedge\neg z\rangle}$${\langle z,\neg z\rangle}$${\langle z,1\rangle}$${\langle z,0\rangle}$${\langle 0,1\rangle}$

### v2

@@ 4.3. A universal-algebraic methodology for the e-generalization type
Assume first that the type of $\mathscr{G}(h)$ is nullary or infinitary; then we prove that there cannot be a finite minimal complete set of solutions to $h$. Assume by way of contradiction that there is, and call it $\mathscr{M}=\{g_{1},\ldots,g_{n}\}$, where each $g_{i}:{\bf F}_{\mathsf{V}}(z)\to{\textbf{P}}_{i}$ for some projective algebra ${\textbf{P}}_{i}\in\mathsf{V}$. Consider now the set of congruences $\{\theta_{i}=\ker(g_{i}):i=1,\ldots,n\}\subseteq\mathscr{G}(h)$; since the type of $\mathscr{G}(h)$ is nullary or infinitary, there must be an exact congruence $\theta\in\mathscr{G}(h)$ such that $\theta\not\subseteq\theta_{i}$ for all $i=1,\ldots,n$. For otherwise one would have that the type of $\mathscr{G}(h)$ is finitary.
@@
For (2), suppose now that the type of $\mathscr{G}(h)$ is finitary. Then the same argument as above, setting $n=1$ and $\mathscr{M}=\{g_{1}\}$, shows that the e-generalization type of $h$ cannot be unitary. ∎
@@
Corollary 4.22.
Every variety whose 1-generated free algebra is trivial has unitary e-generalization type. In particular: lattices, semilattices, all varieties without constants and whose operations are idempotent.
Proof.
In Example 4.14 we have seen that every variety $\mathsf{V}$ whose 1-generated free algebra is trivial has the 1ESP property. This includes lattices, semilattices, and more generally all varieties without constants and whose operations are idempotent. Since the 1-generated free algebra is trivial, the type of $\mathscr{G}(h)$ is clearly unitary for all problems $h$; it follows from Corollary 4.20 that every such variety has unitary $\kappa$-type for every cardinal $\kappa\leq\omega$. ∎
@@
It is proved in [19] that idempotent anti-unification is infinitary, which seems to contradict the result outlined in the above corollary. However, the result in [19] makes key use of two constants, assumed to be in the language, in order to show that the type is (at least) infinitary. If we add constants, the 1-generated free algebra is not trivial, and therefore there is no contradiction. We also note that our result is incomparable with the one in [19], in the sense that while ours is more general with respect to the fact that we do not restrict the equational theory to idempotency (our result holds for any equational theory that includes idempotency of all symbols), we do need to restrict to not having constants.
@@
The previous lemma entails a sufficient condition for varieties to have unitary e-generalization type.
Theorem 4.25.
@@
If $h:{\bf F}_{\mathsf{V}}(z)\to\prod_{k=1}^{m}{\textbf{E}}_{k}$ is any e-generalization problem, under our hypothesis, $\prod_{k=1}^{m}{\textbf{E}}_{k}$ is finitely generated and projective and hence its type is unitary by Lemma 4.24. Thus, the $\kappa$-type of $\mathsf{V}$, for all cardinals $\kappa\leq\omega$, is unitary as well. ∎
The following result closes this section, using the previous theorem to show the unitary type of the varieties of structures considered in the examples throughout this section [^2] Baader in [6] studies the generalization type of what are called commutative theories, which encompass commutative monoids and abelian groups, showing indeed unitary type. However, the generalization order considered in that paper is seemingly different from the one we consider here, as it involves ordering the testifying substitutions and not just the terms..
Corollary 4.26.
Abelian groups, commutative monoids, and commutative semigroups have unitary e-generalization type.
Proof.
In Example 4.8 we have seen that in the varieties of interest the only 1-generated exact algebras are the free 1-generated algebra itself, ${\bf F}_{\mathsf{V}}(z)$ and, in all cases except commutative semigroups, the trivial algebra (which in such cases is projective). Thus, all nontrivial finite direct products of 1-generated exact algebras are finite powers of ${\bf F}_{\mathsf{V}}(z)$.
In all three cases, the $n$-th power ${\bf F}_{\mathsf{V}}(z)^{n}$ is isomorphic to the $n$-generated free algebra. To see this, notice that for all three varieties under consideration the $n$-generated free algebra is the algebra of polynomials (over the appropriate language) written over $n$-variables. Specifically, for abelian groups, commutative monoids, and commutative semigroups one gets $\mathbb{Z}[x_{1},\ldots,x_{n}]$, $\mathbb{N}[x_{1},\ldots,x_{n}]$, and $\mathbb{N}^{+}[x_{1},\ldots,x_{n}]$ respectively. It is then easy to realize that these are isomorphic to, respectively, ${\textbf{Z}}^{n}$, ${\textbf{N}}^{n}$, and $({\textbf{N}}^{+})^{n}$. Therefore, finite direct products of 1-generated exact algebras are (free hence) projective, and we can apply Theorem 4.25 and conclude that all the three varieties above have unitary e-generalization type. ∎
Note that if we drop commutativity and consider groups, monoids, and semigroups, finite direct products of 1-generated exact algebras are not projective and hence one cannot apply Theorem 4.25. Indeed, the 1-generated exact algebras are the same as their commutative versions (see Example 4.8) and hence so are their direct products, which are then commutative. However, these direct products cannot embed into any free algebra since no nontrivial subalgebra of free groups/monoids/semigroups is commutative. Hence finite direct products of 1-generated exact algebras are neither exact nor projective.
@@ 5. Applications to (algebraizable) logics
Given this general perspective, we now investigate some of the most well-known (algebraizable) logics and obtain examples of unitary type for the varieties of Boolean, Gödel, and Kleene algebras, and nullary type for MV-algebras.
@@ 5.1. Classical logic
$\Delta$$\langle z,1\rangle$$\langle\neg z,1\rangle{\nabla}$
Figure 2. ${\bf F}_{\mathsf{BA}}(z)$ and its congruence lattice, where $\langle z,1\rangle$ and $\langle\neg z,1\rangle$ represent, respectively, the congruences generated by $(z,1)$ and $(\neg z,1)$.
@@
Observe that one could have obtained the above result also utilizing Theorem 4.25, since all finitely generated Boolean algebras, and hence in particular finite direct products of 1-generated exact Boolean algebras, are projective.
5.2. Gödel-Dummett logic and locally tabular many-valued logics
@@
Since all non-trivial finite Gödel algebras are projective (see [28, Proposition 2.5] for a proof based on a Priestley-like duality, and [2, Theorem 5.17] for an algebraic proof), $\mathsf{GA}$ has the 1EP property. However, it is not hard to see that the 1ESP fails, as the following example shows.
@@
Nonetheless, we can use Theorem 4.25 to prove that Gödel algebras have unitary e-generalization type. Indeed, all non-trivial finite Gödel algebras are projective, and since Gödel algebras are a locally finite variety, finite direct products of 1-generated exact algebras are finite. We can then conclude the following.
@@
More generally, one can prove uniformly the unitarity of the e-generalization type for all locally finite varieties of (0-bounded) hoops, of which Gödel algebras and Boolean algebras are particular examples. Hoops are a variety of residuated monoids which, particularly in their 0-bounded version, constitute the equivalent algebraic semantics of some of the most relevant many-valued logics. For instance, all logics in Hájek’s framework, among which: Łukasiewicz logic, Gödel Dummett logic, Basic Logic (see [40, 1]). While we preferred to isolate the above cases as a mean to better illustrate our methodology, and because they have a particular interest on their own, let us now present a coincise uniform argument for varieties of hoops.
The key is given by results in [2], specifically Theorem 5.3 and Corollary 5.15, which we summarize in the following statement.
Theorem 5.4 ([2]).
In every locally finite variety of (0-bounded) hoops, finitely generated projective algebras coincide with finite algebras (that have a homomorphism to the 2-element Boolean algebra).
This allows one to apply Theorem 4.25.
Theorem 5.5.
All locally finite varieties of hoops and 0-bounded hoops have unitary e-generalization type.
Proof.
For locally finite variety of hoops, notice that finite direct products of $1$-generated algebras are finite in a locally finite variety, and hence projective by Theorem 5.4. Therefore, one can apply Theorem 4.25 to obtain the unitarity of the type.
For the 0-bounded case, one can apply the same reasoning, modulo the observation that every exact algebra in a variety of 0-bounded hoops has a homomorphism to the 2-element Boolean algebra, and therefore so does any direct product of them. Indeed, every exact algebra embeds into some free algebra by definition, and the 0-generated free algebra is a quotient of any other free algebra; in any nontrivial variety of 0-bounded hoops $\mathsf{V}$, ${\textbf{F}}_{\mathsf{V}}(\emptyset)$ coincides with the 2-element Boolean algebra. This concludes the proof. ∎
Some of the most relevant cases, besides Boolean and Gödel algebras, are summarized in the following corollary.
Corollary 5.6.
The following varieties, and the corresponding many-valued logics, have unitary e-generalization type: all subvarieties of Gödel algebras; all locally finite subvarieties of MV-algebras and BL-algebras; all subvarieties of Gödel hoops; all locally finite subvarieties of Wajsberg hoops and basic hoops.
@@
Let us now consider the 3-valued Kleene logic, also known as Kleene’s “strong logic of indeterminacy”; this is a generalization of classical logic which adds to the intended model of the logic a third indeterminate value to the truth and falsum constants [43]. Its algebraic semantics, the variety $\mathsf{KA}$ of Kleene algebras, is a subvariety of bounded distributive lattices with an involution $\neg$ [42]. In particular $\mathsf{KA}$ is axiomatized by $x\approx\neg\neg x$, De Morgan identity $x\wedge y\approx\neg(\neg x\lor\neg y)$, and Kleene identity $x\wedge\neg x\leq y\vee\neg y$. As it is well-known $\mathsf{KA}$ is generated by its standard model which is the 3-element algebra ${\textbf{K}}_{3}=(\{0,a,1\},\wedge,\vee,\neg,0,1)$ where $0<a<1$ and $\neg 0=1,\neg 1=0,\neg a=a$.
The free $1$-generated algebra ${\bf F}_{\mathsf{KA}}(z)$ is depicted in Figure 3, together with its quotients: ${\bf 4}_{\mathsf{KA}}$ and ${\bf 2}_{\mathsf{KA}}$ (respectively, the 4-element and 2-element Boolean algebras seen as Kleene algebras), the Kleene chains ${\bf K}_{4}$ and ${\bf K}_{3}$, of four and three elements, and the trivial algebra.
[…]
01{1}
01{1}
01{1}
01{1}
@@
(d) $\neg x\leq x,x\land\neg y\leq\neg x\lor y\mbox{ implies }\neg y\leq y.$
We observe that if a Kleene algebra has a negation fixpoint $f=\neg f$ it is not exact, because the above quasi-equation (d) fails. Indeed, $\neg f\leq f$ and $f=f\wedge\neg 0\leq\neg f\vee 0=\neg f$ but $\neg 0=1>0$.
@@
The 1-generated exact Kleene algebras are ${\bf F}_{\mathsf{KA}}(z)$, ${\bf K}_{4}$, and ${\bf 2}_{\mathsf{KA}}$, and they are all projective.
@@
Let us show that, among the 1-generated algebras depicted in Figure 3, only ${\bf F}_{\mathsf{KA}}(z)$, ${\bf 4}_{\mathsf{KA}}$, ${\bf K}_{4}$, ${\bf 2}_{\mathsf{KA}}$ are exact. First, ${\bf F}_{\mathsf{KA}}(z)$ is a retract of itself and it is hence exact and projective. It is also clear that ${\bf 2}_{\mathsf{KA}}$ is a retract of ${\bf F}_{\mathsf{KA}}(z)$ and therefore it is projective and hence exact.
The same argument holds for ${\bf K}_{4}$. Indeed, naming its lattice reduct $0<\neg x<x<1$, consider the embedding into ${\bf F}_{\mathsf{KA}}(z)$ via the map $i$ such that $i(1)=1$, $i(x)=z\vee\neg z$, $i(\neg x)=z\wedge\neg z$, $i(0)=0$. The surjective homomorphism $j:{\bf F}_{\mathsf{KA}}(z)\to{\bf K}_{4}$ defined by $j(z)=j(z\lor\neg z)=x$ and $j(\neg z)=j(z\land\neg z)=\neg x$, composed with $i$ gives the identity on ${\bf K}_{4}$, which is therefore a retract of the free algebra hence projective.
It is left to prove that the other 1-generated algebras are not exact. The trivial algebra is obviously not isomorphic to a subalgebra of any non-trivial algebra and therefore it cannot be exact. As for ${\textbf{4}}_{\mathsf{KA}}$, $1$ is not join irreducible and therefore the claim follows by Lemma 5.7. Finally, ${{\textbf{3}}}_{\mathsf{KA}}$ is not exact since it has a negation fixpoint and, as discussed above, this implies a failure of identity (d) which is necessary for exactness by Lemma 5.7. ∎
@@
(e) $\iota(x)=\bigwedge A-\{\neg a:x\leq_{{\textbf{A}}}a\}.$
@@
[…]